OpenStudy (anonymous):

I really need help! Write an equation of a line parallel to y = 1 over 4x − 6 in slope-intercept form that passes through the point (−1, 5).

OpenStudy (anonymous):

@dan815 can you help please?

Nnesha (nnesha):

what is a slope in this equation ??? y = 1 over 4x - 6

OpenStudy (anonymous):

there is answer choices to one sec!

Nnesha (nnesha):

its oaky we can solve it first don't need to post answer choices :)

Nnesha (nnesha):

okay np y = mx + b this is slope intercept form where m is slope which is always with x and b is y-intercept what is m this equation \[y = \frac{ 1 }{ 4 }x + 6\]

OpenStudy (anonymous):

So we subtract 6?

Nnesha (nnesha):

nope yo don't have to d anything yet i'm asking what is m ???

OpenStudy (anonymous):

Oh 1/4

OpenStudy (anonymous):

the slope

Nnesha (nnesha):

yep right :-) now you need parallel line equation so parallel lines should have same slope slope of new equation its going to be 1/4 :-)

OpenStudy (anonymous):

Okay

Nnesha (nnesha):

\[y = \frac{ 1 }{ 4}x + b\] now use that order pair were first value is x and 2nd one is y ( x , y) and substitute x and y values into the equation and then solve for b

OpenStudy (anonymous):

how do i solve for b?

OpenStudy (anonymous):

Would it be y=1/4x-1

OpenStudy (anonymous):

Answer choices: y=1/4x-1 y=1/4x+5.25 y=-4x+5 y=-4x-1.25

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

wait how ??

Nnesha (nnesha):

 replace x by -1 and y by 5